\begin{thebibliography}{99}

\bibitem{Alk}
Yu.\,A.~Alkhutov,
The Harnack inequality and the H\"{o}lder property of solutions of nonlinear elliptic equations
with a nonstandard growth condition (Russian),
Differ. Uravn. \textbf{33} (1997), no.~12, 1651--1660;
translation in Differential Equations \textbf{33} (1997), no.~12, 1653--1663 (1998).

\bibitem{AlkSur}
Yu. A. Alkhutov, M. D. Surnachev,
H\"{o}lder Continuity and Harnack's Inequality
for $p(x)$-Harmonic Functions, Proceedings of the Steklov Inst. of Math.,  \textbf{308}  (2020),   1--21.




\bibitem{BarColMin1}
P.~Baroni, M.~Colombo, G.~Mingione,
Harnack inequalities for double phase functionals,
Nonlinear Anal. \textbf{121} (2015), 206--222.



\bibitem{BarColMin2}
P.~Baroni, M.~Colombo, G.~Mingione,
Non-autonomous functionals, borderline cases and related
function classes,
St. Petersburg Math. J. \textbf{27} (2016), 347--379.



\bibitem{BarColMin3}
P. Baroni, M. Colombo, G. Mingione,
Regularity for general functionals with double phase, Calc. Var.
Partial Differential Equations \textbf{57}, 62 (2018). %, no. 2, Paper No. , 48 pp.



\bibitem{BenHarHasKar}
A. Benyaiche, P. Harjulehto, P. H\"{a}st\"{o}, A. Karppinen,
The weak Harnack inequality for unbounded supersolutions
of equations with generalized Orlicz growth, J. of Diff. Equations \textbf{275}  (2021), 790--814.



\bibitem{BurSkr}
K.\,O. Buryachenko, I.\,I. Skrypnik,
Local continuity and Harnack's inequality for double-phase parabolic equations,
 Potential Analysis \textbf{56}  (2020), 137--164.



\bibitem{ColMin1}
M.~Colombo, G.~Mingione,
Bounded minimisers of double phase variational integrals,
Arch. Rational Mech. Anal.  \textbf{218} (2015), no. 1, 219--273.



\bibitem{ColMin2}
M.~Colombo, G.~Mingione, Regularity for double phase variational problems,
Arch. Rational Mech. Anal.  \textbf{215} (2015), no. 2, 443--496.



\bibitem{ColMin3}
M. Colombo, G. Mingione,
Calderon-Zygmund estimates and non-uniformly elliptic operators,
J. Funct. Anal. \textbf{270} (2016), 1416--1478.

\bibitem{CupMarMas}
G. Cupini, P. Marcellini, E. Mascolo,
Local boundedness of minimizers with limit growth conditions,
J. Optim. Theory Appl. \textbf{166} (2015), 1--22.


\bibitem{DiBGiaVes}
E. DiBenedetto, U. Gianazza, V. Vespri, Local clustering of the non-zero set of functions in $W^{1,1}(E)$,
Atti Accad. Naz. Lincei Rend. Lincei Mat. Appl. \textbf{17}(3) (2006) , 223--225.


\bibitem{DiBTru}
E. DiBenedetto, N.S. Trudinger, Harnack inequalities for quasi-minima of variational integrals, Ann.
Inst. Henri Poincare, Analyse Non Lineaire  \textbf{1}(4)  (1984) ,  295--308


\bibitem{HadSkrVoi}
O. V. Hadzhy, I. I. Skrypnik, M. V. Voitovych, Interior continuity, continuity up to the
boundary and Harnack's inequality for double-phase elliptic equations with non-logarithmic
growth, Math. Nachrichten, in press.

\bibitem{HarKinLuk}
P. Harjulehto, J. Kinnunen, T. Lukkari,
Unbounded supersolutions of nonlinear equations with nonstandard growth,
Bound. Value Probl. 2007 (2007) 48348

\bibitem{HarHasLee}
P. Harjulehto, P. H\"{a}st\"{o}, M. Lee,
H\"{o}lder continuity of quasiminimizers and $\omega$-minimizers of functionals with generalized
Orlicz growth,  Ann. Sc. Norm. Super Pisa Cl. Sci \textbf{5}  XXII
(2021), no.2, 549--582.

\bibitem{HarHasToi}
P. Harjulehto, P. H\"{a}st\"{o}, O. Toivanen,
H\"{o}lder regularity of quasiminimizers under generalized growth conditions,
Calc. Var. Partial Differential Equations \textbf{56} (2017), no. 2, Art. 22, 26 pp.



\bibitem{KrySaf}
N.\,V. Krylov, M.\,V. Safonov,
A property of the solutions of parabolic equations with measurable coefficients,
Izv. Akad. Nauk SSSR Ser. Mat. \textbf{44} (1980), no. 1, 161--175 (in Russian).



\bibitem{LadUra}
O.\,A.~Ladyzhenskaya, N.\,N.~Ural'tseva,
Linear and Quasilinear Elliptic Equations,
Nauka, Moscow, 1973.


\bibitem{Lie}
G.\,M.~Lieberman,
The natural generalization of the natural conditions of
Ladyzhenskaya and Ural'tseva for elliptic equations,
Comm. Partial Differential Equations \textbf{16} (1991), no. 2-3, 311--361.

\bibitem{LisSkr}
V. Liskevich, I. I. Skrypnik, Harnack inequality and continuity of solutions to elliptic equations with nonstandard
growth conditions and lower orfer terms, Ann. Mat. Pura ed Appl. \textbf{189} (2010), 335--356.

\bibitem{MizOhnShi}
Y. Mizuta, T. Ohno, T. Shimomura,
Sobolev's theorem for double phase functionals,
Math. Inequal. Appl.  \textbf{23} (2020), no. 1,  17--33.

\bibitem{Ok}
J. Ok,
Regularity for double phase problems under additional integrability assumptions,
Nonlinear Anal. \textbf{194} (2020) 111408.

\bibitem{RagTac}
M. A. Ragusa, A. Tachikawa,
Regularity for minimizers for functionals of double phase with variable exponents,
Adv. Nonl. Anal. \textbf{9}(1) (2020), 710--728.

\bibitem{SavSkrYev}
M. O. Savchenko, I. I. Skrypnik, Y. A. Yevgenieva,
Continuity and Harnack inequalities for local minimizers of non uniformly elliptic functionals with generalized Orlicz growth
under the non-logarithmic conditions, Nonl. Analysis \textbf{230} (2023), 113221.

\bibitem{ShaSkrVoi}
M.\,A. Shan,  I.\,I. Skrypnik,  M.\,V. Voitovych,
Harnack's inequality for quasilinear elliptic equations with
generalized Orlicz growth, Electr. J. of Diff. Equations, \textbf{2021} (2021), no. 27,  1--16.




\bibitem{SkrVoi1}
I.\,I. Skrypnik, M.\,V. Voitovych,
$\mathfrak{B}_{1}$ classes of De Giorgi, Ladyzhenskaya and Ural'tseva
and their application to elliptic and parabolic equations with nonstandard growth,
Ukr. Mat. Visn. \textbf{16} (2019), no. 3, 403--447.



\bibitem{SkrVoi2}
I.\,I. Skrypnik, M.\,V. Voitovych,
$\mathcal{B}_{1}$ classes of De Giorgi-Ladyzhenskaya-Ural'tseva and
their applications to elliptic and parabolic equations with generalized Orlicz
growth conditions, Nonlinear Anal. \textbf{202} (2021) 112--135.



\bibitem{Sur}
M. D. Surnachev, On the weak Harnack inequality for the parabolic $p(x)$- Laplacian,
Asymptotic Analysis, DOI:10.3233/ASY-211746 (2021).



\end{thebibliography}

\documentclass[a4paper,11pt,fullpage]{article}
\usepackage[english]{babel}

\usepackage[left=1in,right=1in,top=1in,bottom=1in]{geometry}

\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{amsxtra}
\usepackage{latexsym}
\usepackage{ifthen}
\usepackage{cite}
\usepackage{esint}
\usepackage[cp1250]{inputenc}

\usepackage{epic}
\usepackage{eepic}

\DeclareMathOperator*{\osc}{osc}

\numberwithin{equation}{section}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}{Lemma}[section]
\newtheorem{remark}{Remark}[section]

\def\XXint#1#2#3{{\setbox0=\hbox{$#1{#2#3}{\int}$}
     \vcenter{\hbox{$#2#3$}}\kern-.5\wd0}}

\begin{document}

\title{A note on the weak Harnack inequality for unbounded  minimizers of elliptic functionals with generalized Orlicz growth
}

\author{Mariia O. Savchenko, Igor I. Skrypnik, Yevgeniia A. Yevgenieva
 }

  \maketitle

  \begin{abstract} We prove the weak Harnack inequality for the functions $u$ which belong to the corresponding De Giorgi classes $DG^{-}(\Omega)$ under the additional assumption that $u\in L^{s}_{loc}(\Omega)$ with some  $s> 0$.  In particular, our result covers new cases of  functionals with a variable exponent or double-phase functionals under the non-logarithmic condition.

\textbf{Keywords:}
non-autonomous functionals, unbounded minimizers, weak Harnack inequality.

\textbf{MSC (2010)}: 35B40, 35B45, 35B65.

%\textbf{Declarations of interest}: none.

\end{abstract}

\pagestyle{myheadings} \thispagestyle{plain}
\markboth{Mariia O. Savchenko, Igor I. Skrypnik, Yevgeniia A.Yevgenieva}
{A note on the weak Harnack inequality....}

\section{Introduction and Main Results}\label{Introduction}
It is known that for integrands with $p,q$-growth, it is crucial that the gap between $p$ and $q$ is not too large. Otherwise, in the case
$q> \dfrac{np}{n-p}$, $p<n$ there exist unbounded minimizers and at the same time, the constant in Harnack inequality cannot be independent of the function, in contrast to the standard case, i.e. if $p=q$ (we refer the reader to \cite{Alk, AlkSur, BarColMin1, BarColMin2, BarColMin3, BenHarHasKar, BurSkr, ColMin1, ColMin2, ColMin3, CupMarMas, HadSkrVoi, HarKinLuk, HarHasLee, HarHasToi, Lie, LisSkr, MizOhnShi, Ok, RagTac, SavSkrYev, SkrVoi1, SkrVoi2, Sur} for results, references, historical notes and extensive survey of regularity issues). It was Ok \cite{Ok}, who proved the boundedness of minimizers of elliptic functionals of double-phase type in the case $q>\dfrac{np}{n-p}$ under some additional assumption. More precisely, under the condition $\osc\limits_{B_{r}(x_{0})}a(x)\leqslant A r^{a}$, $a\geqslant q-p$ the function $u$ belonging to the corresponding De Giorgi class $DG^{+}(\Omega)$ is bounded by a constant depending on $||u||_{L^{s}}$ with $s \geqslant \dfrac{(q-p) n}{a+p-q}$. This condition, for example, gives a possibility to improve the regularity results \cite{BarColMin1, BarColMin2, BarColMin3, ColMin1, ColMin2} for unbounded minimizers with constant depending on $||u||_{L^{s}}$. The weak Harnack inequality for unbounded supersolutions of the corresponding elliptic equations with generalized Orlicz growth  under the so-called logarithmic conditions was proved by the Moser method  in \cite{BenHarHasKar}.

It seems that for the corresponding De Giorgi classes $DG^{-}(\Omega)$ this question remains open even under the so-called logarithmic conditions, i.e. if $\lambda(r) \equiv 1$ (see condition ($\varPhi_{\lambda}$) below). In this note, we will prove the weak Harnack inequality for functions belonging to the corresponding  elliptic De Giorgi classes $DG^{-}(\Omega).$

We write  $W^{1,\varPhi(\cdot)}(\Omega)$ for the class of functions $u\in W^{1,1}(\Omega)$ with
$\int\limits_{\Omega}\varPhi(x, |\nabla u|)dx < \infty$ and we say that a measurable function $u:\Omega\rightarrow \mathbb{R}$ belongs to the elliptic class $DG^{\pm}_{\varPhi}(\Omega)$ if $u\in W^{1,\varPhi(\cdot)}(\Omega)$ and there exist numbers  $c >0$, $q>1$ such that for any ball $B_{8r}(x_{0})\subset\Omega$, any $k \in \mathbb{R}$  and any $ \sigma\in(0,1)$  the following inequalities hold:

\begin{equation}\label{eq1.1}
\int\limits_{A^{\pm}_{k,r(1-\sigma)}}  \varPhi\big(x,|\nabla u|\big) dx \leqslant  \frac{c}{\sigma^{q}}\,\int\limits_{A^{\pm}_{k,r}}  \varPhi\bigg(x,\frac{(u-k)_{\pm}}{r}\bigg) dx,
\end{equation}
here  $(u-k)_{\pm}:=\max\{\pm(u-k), 0\}$, $A^{\pm}_{k,r}:=B_{r}(x_{0})\cap \{(u-k)_{\pm}>0\}$.

Further, we suppose that $\varPhi(x, v):\Omega\times \mathbb{R}_{+}\rightarrow \mathbb{R}_{+}$ is a non-negative function satisfying the following properties: for any $x\in\Omega$
 the function $ v\rightarrow \varPhi(x, v)$ is increasing and
 $\lim\limits_{ v\rightarrow0}\varPhi(x, v)=0$,
 $\lim\limits_{ v\rightarrow +\infty}\varPhi(x, v)=+\infty$. We also assume that

\begin{itemize}
\item[($\varPhi$)]
There exist $1<p<q$ such that
for $x\in \Omega$ and for $ w\geqslant v > 0$ there holds
\begin{equation*}
\left( \frac{w}{v} \right)^{p} \leqslant\frac{\varPhi(x, w)}{\varPhi(x, v)}\leqslant
 \left( \frac{w}{v} \right)^{q}.
\end{equation*}
\end{itemize}

\begin{itemize}
\item[($\varPhi_{\lambda}$)]
There exist $s >0$, $R >0$ and  continuous, non-decreasing function
$\lambda(r) \in(0,1)$ on the interval $(0,R)$, $\lim\limits_{r \rightarrow 0} \lambda(r)=0$, $\lim\limits_{r \rightarrow 0} \dfrac{r}{\lambda(r)}=0$,   such that for any $B_{r}(x_{0}) \subset B_{R}(x_{0}) \subset \Omega$ and some
$A >0 $ there holds
$$
\varPhi^{+}_{B_{r}(x_{0})}\bigg( \frac{\lambda(r) v}{r^{1+\frac{n}{s}}}\bigg)\leqslant A\,\, \varPhi^{-}_{B_{r}(x_{0})}\bigg(\frac{\lambda(r) v}{r^{1+\frac{n}{s}}}\bigg),\quad r^{1+\frac{n}{s}} \leqslant \lambda(r) v\leqslant 1,
$$
here $\varPhi^{+}_{B_{r}(x_{0})}( v):=\sup\limits_{x\in B_{r}(x_{0})}\varPhi(x, v),\quad  \varPhi^{-}_{B_{r}(x_{0})}( v):=\inf\limits_{x\in B_{r}(x_{0})}\varPhi(x, v),\quad v > 0 $.
\end{itemize}
For the function $\lambda(r)$ we also need the following condition
\begin{itemize}
\item[($\lambda$)]
For any $0< r < \rho< R$ there holds
\begin{equation*}
\lambda(r) \geqslant \lambda(\rho) \bigg(\frac{r}{\rho}\bigg)^{b},
\end{equation*}
with some $b \geqslant 0$.
\end{itemize}
For the function $\lambda(r)=\left[\log\dfrac{1}{r}\right]^{-\frac{\beta}{q-p}}$, $\beta \geqslant 0$ this condition holds evidently, provided that $R$ is small enough.
\begin{remark}\label{rem1.1}
Consider the function $\varPhi(x, v):= v^{p}+a(x)v^{q}$, $a(x) \geqslant 0$, $\osc\limits_{B_{r}(x_{0})}a(x) \leqslant K r^{a} \big[\log\frac{1}{r}\big]^{\beta}$, $a\in (0,1]$, $\beta \geqslant 0$, $K>0$. Evidently condition ($\varPhi_{\lambda}$) holds with $\dfrac{n(q-p)}{a+p-q} \leqslant s \leqslant \infty$, $a\geqslant q-p$, $\lambda(r):= \big[\log\frac{1}{r}\big]^{-\frac{\beta}{q-p}}$ and $A=K^{q-p}$.

For the function $\varPhi(x, v):=v^{p(x)}$, $\osc\limits_{B_{r}(x_{0})}p(x)\leqslant\dfrac{L}{\log\frac{1}{r}}$, $L>0$ condition ($\varPhi_{\lambda}$) holds with $s>0$, $\lambda(r)\equiv 1$ and $A=\exp\big(L(1+\frac{n}{s})\big)$.
\end{remark}
\begin{remark}\label{rem1.2}
We note that conditions ($\varPhi_{\lambda}$) and ($A1-s_{*}$) with $s_{*}=\dfrac{ns}{n+s}$ from \cite{BenHarHasKar} essentially coincide in the case $\lambda(r)\equiv 1$.
\end{remark}

We refer to the parameters $n$, $p$, $q$, $s$,  $c$, $A$   as our structural  data, and we write $\gamma$ if it can be quantitatively determined a priory in terms of the above quantities. The generic constant $\gamma$ may change from line to line.

Our main result reads as follows.
\begin{theorem}\label{th1.1}
Let $u\in DG^{-}(\Omega)$, $u\geqslant 0$, let conditions ($\varPhi$), ($\varPhi_{\lambda}$), ($\lambda$) be fulfilled. Let $B_{8\rho}(x_{0})\subset B_{R}(x_{0})\subset \Omega$, let additionally
$u \in L^{s}_{loc}(\Omega)$ with some $s\geqslant q-p$ and $\left(\int\limits_{B_{2\rho}(x_{0})} u^{s} \right)^{\frac{1}{s}} \leqslant d$. Then there exists a positive constant $C$ depending only
on the known parameters and $d$, such that
\begin{equation}\label{eq1.2}
\left(\,\,\fint\limits_{B_{\rho}(x_{0})} (u+\rho )^{\theta} dx\right)^{\frac{1}{\theta}}\leqslant \frac{C}{\lambda(\rho)} \left(\inf\limits_{B_{\frac{\rho}{2}}(x_{0})} u +\rho \right),
\end{equation}
where $\fint\limits_{B_{\rho}(x_{0})}u^{\theta} dx:= |B_{\rho}(x_{0})|^{-1} \int\limits_{B_{\rho}(x_{0})}u^{\theta} dx$ and $\theta >0$ is some fixed number  depending only on the known data.
\end{theorem}

The conditions of Theorem~\ref{th1.1} are precise, we refer the reader to \cite{BenHarHasKar} for the examples.
In the case  $s=\infty$, Theorem \ref{th1.1}  was proved in \cite{BarColMin1, SavSkrYev}.

The main difficulty arising in the proof of our main result, Theorem \ref{th1.1}, is related to the so-called theorem on the expansion of positivity. Roughly speaking, having information on the measure of the ''positivity set'' of $u$ over the ball $B_{r}(\bar{x})$:
$$|\{x \in B_{r}(\bar{x}) : u(x) \geqslant N \}| \geqslant \alpha|B_{r}(\bar{x})|,$$
with some $r, N >0$ and  $\alpha \in (0, 1)$, we need to  translate it into the expansion of the set of positivity  to a ball $B_{2r}(\bar{x})$. Difficulties arise not only due to the presence of a factor $\lambda(r) $ in condition ($\varPhi_{\lambda}$), but also due to the presence of  the second term on the right-hand side of inequality \eqref{eq1.1}.  We do not use the classical covering argument of  Krylov and Safonov \cite{KrySaf}, DiBenedetto and  Trudinger \cite{DiBTru} as it was done in the ''bounded'' case, i.e. if $s= + \infty$ (see e.g. \cite{BarColMin1}), instead we use the  local clustering lemma due to DiBenedetto, Gianazza, and Vespri \cite{DiBGiaVes} and moreover, instead of $\sup\limits_{B_{2r}(\bar{x})}u$ we are forced to use averages of $u$ over the ball $B_{2r}(\bar{x})$.

The rest of the paper contains proof of the above theorem. In Section \ref{Sec2} we collect some auxiliary propositions and required integral estimates of functions belonging to the corresponding De Giorgi classes. Section \ref{Sec3} contains the proof of weak Harnack inequality, Theorem~\ref{th1.1}.
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\section{Auxiliary Material and Integral Estimates}\label{Sec2}

\subsection{ Local Clustering Lemma}\label{subsect2.1}

The following  lemma will be used in the sequel, it is  the local clustering lemma, see \cite{DiBGiaVes}.

\begin{lemma}\label{lem2.1}
Let $K_{r}(y)$ be a cube in $\mathbb{R}^{n}$ of edge $r$ centered at $y$ and let $u\in W^{1,1}(K_{r}(y))$ satisfy
\begin{equation}\label{eq2.1}
||(u-k)_{-}||_{W^{1,1}(K_{r}(y))} \leqslant \mathcal{K}\,k\,r^{n-1},\,\,\,\,\,\,and\,\,\,\,\,\, |\{K_{r}(y) : u\geqslant k \}|\geqslant \alpha |K_{r}(y)|,
\end{equation}
with some $\alpha\in (0,1)$, $k\in\mathbb{R}^{1}$ and $\mathcal{K} >0$. Then for any $\xi \in (0,1)$ and any $\nu\in (0,1)$ there exists $\bar{x} \in K_{r}(y)$ and $\delta=\delta(n) \in (0,1)$ such that
\begin{equation}\label{eq2.2}
|\{K_{\bar{r}}(\bar{x}):  u\geqslant \xi\,k \}| \geqslant (1-\nu) |K_{\bar{r}}(y)|,\,\,\, \bar{r}:=\delta \alpha^{2}\frac{(1-\xi)\nu}{\mathcal{K}}\,r.
\end{equation}
\end{lemma}

\subsection{Local Energy Estimates}\label{subsec2.2}

The following lemma is a consequence of inequalities \eqref{eq1.1}.
\begin{lemma}\label{lem2.2}
Let $u\in DG^{-}(\Omega$), $u\geqslant 0$,  $B_{r}(\bar{x})\subset B_{\rho}(x_{0}) \subset B_{8\rho}(x_{0})  \subset \Omega$, and let condition ($\varPhi$) holds, then for  any $k> 0$, any $\sigma \in(0,1)$ there holds
\begin{equation}\label{eq2.3}
\int\limits_{B_{r(1-\sigma)}(\bar{x})} | \nabla (u-\lambda(r) k)_{-}|^{p} \, dx \leqslant \gamma \sigma^{-q}
\frac{\varPhi^{+}_{B_{r}(\bar{x})}\big( \frac{\lambda(r) k}{r}\big)}{\varPhi^{-}_{B_{r}(\bar{x})}\big( \frac{\lambda(r) k}{r}\big)} \bigg(\frac{\lambda(r)k}{r}\bigg)^{p} |A^{-}_{\lambda(r)k,r}|.
\end{equation}

If additionally condition ($\varPhi_\lambda$) holds and
\begin{equation}\label{eq2.4}
r \leqslant \lambda(r)  k \leqslant \frac{1}{r^{\frac{n}{s}}},
\end{equation}
then
\begin{equation}\label{eq2.5}
\int\limits_{B_{r(1-\sigma)}(\bar{x})} | \nabla (u-\lambda(r) k)_{-}|^{p} \, dx \leqslant \gamma \sigma^{-q}
 \bigg(\frac{\lambda(r)k}{r}\bigg)^{p} |A^{-}_{\lambda(r)k,r}|.
\end{equation}
\end{lemma}
\begin{proof}
First, note the following Young's inequality
\begin{equation*}
\varPhi_{p}(x,a) b^{p} \leqslant \varPhi(x,a) + \varPhi(x,b),\quad a,b >0, \quad \varPhi_{p}(x,a):=a^{-p}\varPhi(x,a),
\end{equation*}
indeed,  if $b\leqslant a$, then $\varPhi_{p}(x,a) b^{p} \leqslant \varPhi(x,a)$ and if $b\geqslant a$, using the fact that by condition ($\varPhi$), function $\varPhi_{p}(x,a)$ is increasing, we obtain  $\varPhi_{p}(x,a) b^{p} \leqslant \varPhi(x,b)$.

Using this Young's inequality and inequalities \eqref{eq1.1} we get
\begin{multline*}
\int\limits_{B_{r(1-\sigma)}(\bar{x})} \varPhi^{-}_{B_{r}(\bar{x})}\bigg(\frac{\lambda(r) k}{r}\bigg)| \nabla (u-\lambda(r) k)_{-}|^{p} \, dx
\leqslant\\ \leqslant \bigg(\frac{\lambda(r) k}{r}\bigg)^{p}\int\limits_{B_{r(1-\sigma)}(\bar{x})}\varPhi_{p}\bigg(x,\frac{\lambda(r) k}{r}\bigg)|\nabla (u-\lambda(r)k)_{-}|^{p} dx \leqslant \\ \leqslant
\bigg(\frac{\lambda(r) k}{r}\bigg)^{p}\bigg\{\varPhi^{+}_{B_{r}(\bar{x})}\bigg(\frac{\lambda(r) k}{r}\bigg)|A^{-}_{\lambda(r)k,r}|+
\int\limits_{B_{r(1-\sigma)}(\bar{x})}\varPhi\big(x,|\nabla (u-\lambda(r)k)_{-}|\big) dx\bigg\} \leqslant \\ \leqslant
\gamma \sigma^{-q}\bigg(\frac{\lambda(r) k}{r}\bigg)^{p}\varPhi^{+}_{B_{r}(\bar{x})}\bigg(\frac{\lambda(r) k}{r}\bigg)|A^{-}_{\lambda(r)k,r}|,
\end{multline*}
which proves \eqref{eq2.3}. To prove \eqref{eq2.5} we note that by condition ($\varPhi_{\lambda}$)
\begin{equation}\label{eq2.6}
\frac{\varPhi^{+}_{B_{r}(\bar{x})}\big( \frac{\lambda(r) k}{r}\big)}{\varPhi^{-}_{B_{r}(\bar{x})}\big( \frac{\lambda(r) k}{r}\big)}
\leqslant \gamma,
\end{equation}
provided that $r^{1+\frac{n}{s}} \leqslant \lambda(r) r^{\frac{n}{s}}  k \leqslant 1$, which proves \eqref{eq2.5}. This completes the proof of the lemma.
\end{proof}

\subsection{Expansion of the Positivity }\label{subsec2.3}
The following lemma is a consequence of Lemma \ref{lem2.2} and Lemmas $6.2$, $6.3$ from \cite[Chap.\,2]{LadUra}.

\begin{lemma}\label{lem2.3}
Let $u\in DG^{-}(\Omega$), $u\geqslant 0$,  $B_{r}(\bar{x})\subset B_{\rho}(x_{0}) \subset B_{8\rho}(x_{0})  \subset \Omega$,
assume that the number $k>0$ satisfies the condition
\begin{equation}\label{eq2.7}
 \lambda(r)  k \leqslant \frac{1}{r^{\frac{n}{s}}},
\end{equation}
and assume also that with some $\alpha_{0} \in(0,1)$ there holds
\begin{equation}\label{eq2.8}
|\big\{B_{\frac{r}{2}}(\bar{x}): u \geqslant \lambda(r) k\big\}| \geqslant \alpha_{0} |B_{\frac{r}{2}}(\bar{x})| ,
\end{equation}
then there exists number $\eta_{0}\in(0,1)$, depending only on the data and $\alpha_{0}$, such that either
\begin{equation}\label{eq2.9}
\lambda(r) k \leqslant \frac{r}{\eta_{0} },
\end{equation}
or
\begin{equation}\label{eq2.10}
u(x) \geqslant \eta_{0}\,\lambda(r)\,k,\quad x \in B_{r}(\bar{x}).
\end{equation}

\end{lemma}
The proof of the lemma is almost standard. If \eqref{eq2.9} is violated, then by \eqref{eq2.7} inequalities  \eqref{eq2.4} hold. So, \eqref{eq2.5} define the standard De Giorgi classes $DG^{-}_{p}(\Omega)$ with the appropriate choice of number $k$. We refer the reader for the details to Lemmas $6.2$ and $6.3$ of \cite[Chap.\,2]{LadUra}.

\section{Weak  Harnack Inequality, Proof of Theorem \ref{th1.1}}\label{Sec3}

First, we prove the following lemma.
\begin{lemma}\label{lem3.1}
Let $B_{\rho}(x_{0}) \subset B_{8\rho}(x_{0}) \subset \Omega$, $\left(\int\limits_{B_{2\rho}(x_{0})} u^{s} \right)^{\frac{1}{s}} \leqslant d$ and let the following inequality holds
\begin{equation}\label{eq3.1}
\left|\left\{\, B_{\frac{\rho}{2}}(x_{0}): u \geqslant N \,\right\}\right| \geqslant \alpha \left|B_{\frac{\rho}{2}}(x_{0})\right|,
\end{equation}
for some $N >0$ and some  $\alpha \in(0,1)$. Then, under the conditions of Lemma~\ref{lem2.2}, there exist $C_{2}$, $\tau >0$ depending only on the data and $d$ such that either
\begin{equation}\label{eq3.2}
\alpha^{\tau}\leqslant  \frac{C_{2} \rho}{N \lambda(\rho)},
\end{equation}
or
\begin{equation}\label{eq3.3}
\alpha^{\tau}\leqslant \frac{C_{2}}{N \lambda(\rho)} \inf\limits_{B_{\frac{\rho}{2}}(x_{0})} u.
\end{equation}
\end{lemma}
\begin{proof}
Further we will assume that inequality \eqref{eq3.2} is violated, i.e. with some $\tau >0$
\begin{equation}\label{eq3.4}
\alpha^{\tau} N \lambda(\rho) \geqslant  C_{2} \rho.
\end{equation}
Let $\varepsilon\in(0,1)$ be some number to be chosen later. Applying inequality \eqref{eq2.3}  for \\ $(u- \lambda(\rho)\varepsilon N)_{-}$ over the pair of balls $B_{\frac{\rho}{2}}(x_{0})$ and $B_{\rho}(x_{0})$, we obtain
\begin{equation}\label{eq3.5}
\fint\limits_{B_{\frac{\rho}{2}}(x_{0})} |\nabla (u- \lambda(\rho)\varepsilon N)_{-}|^{p}\,dx \leqslant \gamma \,\frac{\varPhi^{+}_{B_{\rho}(x_{0})}\big( \frac{\lambda(\rho)\varepsilon N}{\rho}\big)}{\varPhi^{-}_{B_{\rho}(x_{0})}\big( \frac{\lambda(\bar{\rho})\varepsilon N }{\rho}\big)}\bigg(\lambda(\rho)\frac{\varepsilon N}{\rho}\bigg)^{p}  .
\end{equation}
Now we need to estimate the right-hand side of the last inequality, for this we note that inequality \eqref{eq3.1} yields
\begin{equation}\label{eq3.6}
\left(\fint\limits_{B_{2\rho}(x_{0})} u^{s}\,dx\right)^{\frac{1}{s}}\geqslant  4^{-\frac{n}{s}} \alpha^{\frac{1}{s}} N,
\end{equation}
and moreover
\begin{equation*}
|\{\, B_{\frac\rho2}(x_{0}) : u \geqslant \lambda(\rho)(1+d)^{-1} 4^{-\frac{n}{s}}\alpha^{\frac{1}{s}} N \,\}|\geqslant |\{\, B_{\frac\rho2}(x_{0}) : u \geqslant  N \,\}|\geqslant \alpha |B_{\frac\rho2}(x_{0})|.
\end{equation*}
Choosing $\varepsilon =  (1+d)^{-1} 4^{-\frac{n}{s}} \alpha^{\frac{1}{s}}$, by \eqref{eq3.4}, \eqref{eq3.6}, we obtain
\begin{equation}\label{eq3.7}
\rho \leqslant \lambda(\rho) \varepsilon N \leqslant (1+d)^{-1} \left(\fint\limits_{B_{2\rho}(x_{0})} u^{s}\,dx\right)^{\frac{1}{s}}
\leqslant \frac{d}{(d+1)\rho^{\frac{n}{s}}}\leqslant  \frac{1}{\rho^{\frac{n}{s}}},
\end{equation}
provided that $\tau \geqslant \dfrac{1}{s}$ and $C_{2} \geqslant (1+d) 4^{\frac{n}{s}}$.  Therefore Lemma \ref{lem2.2} and inequalities \eqref{eq3.5}, \eqref{eq3.7} yield
\begin{equation}\label{eq3.8}
\fint\limits_{B_{\frac{\rho}{2}}(x_{0})} |\nabla (u- \lambda(\rho)\varepsilon N)_{-}|^{p}\,dx \leqslant \gamma \,\bigg(\lambda(\rho)\frac{\varepsilon N}{\rho}\bigg)^{p}.
\end{equation}

The local clustering Lemma \ref{lem2.1} with $k=\lambda(\rho)\varepsilon  N$, $\nu=\frac{1}{4}$, $\xi=\frac{1}{4}$, $\mathcal{K}=\gamma $, $r=\frac{\rho}{2}$  implies the existence of a point $\bar{x}\in B_{\frac{\rho}{2}}(x_{0})$ and $\delta \in(0,1)$ depending only on the data, such that
\begin{equation*}
\left|\left\{B_{\bar{r}}(\bar{x}): u \geqslant \frac{\lambda(\bar{r})}{4} \varepsilon N \right\}\right|\geqslant \left|\left\{B_{\bar{r}}(\bar{x}): u \geqslant \frac{\lambda(\rho)}{4} \varepsilon N \right\}\right| \geqslant \frac{3}{4} |B_{\bar{r}}(\bar{x})|,\quad \bar{r}= \delta_{0} \alpha^{2} \rho, \quad \delta_{0}:= \frac{3\delta}{16\mathcal{K}}.
\end{equation*}
Since $\lambda(r)$ is non-decreasing, inequality \eqref{eq3.7} implies
\begin{equation}\label{eq3.9}
\frac{\lambda(\bar{r})}{4}\varepsilon N \leqslant \frac{\lambda(\rho)}{4}\varepsilon N \leqslant \frac{1}{\rho^{\frac{n}{s}}} \leqslant
\frac{1}{\bar{r}^{\frac{n}{s}}},
\end{equation}
and moreover, by condition ($\lambda$) and by \eqref{eq3.4}
\begin{equation*}
\frac{\lambda(\bar{r})}{4}\varepsilon N \geqslant \frac{\lambda(\rho)}{4}\bigg(\frac{\bar{r}}{\rho}\bigg)^{b}\varepsilon N
=\lambda(\rho)\frac{\delta_{0}^{b}}{4^{1+\frac{n}{s}}}(1+d)^{-1} \alpha^{\frac{1}{s}+ 2b} N \geqslant \rho\geqslant \bar{r},
\end{equation*}
provided that $\tau \geqslant 2b +\dfrac{1}{s}$ and $C_{2}\geqslant (1+d) \dfrac{4^{1+\frac{n}{s}}}{\delta_{0}^{b}}$.

So, Lemma \ref{lem2.3} is applicable with $\alpha_{0}=\dfrac{3}{4}$ and $k=\dfrac{\varepsilon}{4} N$, so we obtain with some $\eta_{0} \in(0,1)$ depending only on the data
$$
u(x)\geqslant \frac{\eta_{0}\lambda(\bar{r})\varepsilon N}{4}, \quad x \in B_{2\bar{r}}(\bar{x}),
$$
provided that $C_{2}\geqslant (1+d)\dfrac{4^{1+\frac{n}{s}}}{\eta_{0}\delta_{0}^{b}}$.

Repeating this procedure $j$ times we obtain
\begin{equation}\label{eq3.10}
 u(x)\geqslant \frac{\eta_{0}^{j} \,\lambda(\bar{r})\varepsilon N}{4}, \quad x \in B_{2^{j}\bar{r}}(\bar{x}),
\end{equation}
provided that
\begin{equation}\label{eq3.11}
2^{j} \bar{r}\leqslant \frac{\eta_{0}^{j} \,\lambda(\bar{r})\varepsilon N}{4}\leqslant \dfrac{1}{(2^{j}\bar{r})^{\frac{n}{s}}}.
\end{equation}
 Choose $j$ by the condition $2^{j}\bar{r}=\rho$, that is $2^{j} \delta_{0} \alpha^{2}=1$, the second inequality in \eqref{eq3.11} holds by \eqref{eq3.9}. By \eqref{eq3.4} and  condition ($\lambda$), we have
\begin{equation*}
\eta_{0}^{j} \lambda(\bar{r}) \varepsilon N \geqslant \frac{\delta_{0}^{\log\frac{1}{\eta_{0}}+b}}{(1+d) 4^{\frac{n}{s}}}\,\alpha^{2\log\frac{1}{\eta_{0}}+ 2b+\frac{1}{s}}\,N\,\lambda(\rho) \geqslant \rho= 2^{j} \bar{r},
\end{equation*}
provided that  $\tau=2 \log\dfrac{1}{\eta_{0}}  + 2b +\dfrac{1}{s}$  and $C_{2}=C_{2}(d,\eta_{0},\delta_{0})>0$ is large enough.  Therefore,
inequality \eqref{eq3.10} yields
\begin{equation*}
u(x)\geqslant \frac{\lambda(\rho)}{C_{2}} \alpha^{\tau} N, \quad  x \in B_{\frac{\rho}{2}}(x_{0}),
\end{equation*}
which completes the proof of the lemma.
\end{proof}
To complete the proof of the weak Harnack inequality,  we set  $\bar{m}(\rho)=\dfrac{1}{\lambda(\rho)}\left( \inf\limits_{B_{\frac{\rho}{2}}(x_{0})}u(x)+\rho \right),$ then Lemma \ref{lem3.1}  with  $\theta\in\big(0, \frac{1}{2\tau}\big]$  yields
\begin{multline*}
\fint\limits_{B_{\rho}(x_{0})}u^{\theta}\,dx=\frac{\theta}{|B_{\rho}(x_{0})|}\,\int\limits_{0}^{\infty}|\{B_{\rho}(x_{0}): u(x) >N\}|\,N^{\theta-1}\,dN \leqslant \\ \leqslant  [\bar{m}(\rho)]^{\theta}+ \gamma [\bar{m}(\rho)]^{\frac{1}{\tau}}\int\limits_{\bar{m}(\rho)}^{\infty} N^{\theta-\frac{1}{\tau}-1} \,dN \leqslant \gamma [\bar{m}(\rho)]^{\theta},
\end{multline*}
which proves Theorem \ref{th1.1}.

\vskip3.5mm
{\bf Acknowledgements.} The authors are partially supported due to the project "Mathematical \,\,modeling\,\, of\, complex\,\, dynamical\,\, systems\,\, and\,\, processes\,\, caused\,\, by\,\, the\,\, state\,\, security" (Reg. No. 0123U100853), by the grant of Ministry of Education and Science of Ukraine (Reg. No. 0121U109525) and by the Grant EFDS-FL2-08 of the found The European Federation of Academies of Sciences and Humanities (ALLEA).

%\bigskip
\newpage
CONTACT INFORMATION

\medskip
\textbf{Mariia O.~Savchenko}\\Institute of Applied Mathematics and Mechanics,
National Academy of Sciences of Ukraine, \\ \indent Heneral Batiuk Str. 19, 84116 Sloviansk, Ukraine\\
shan\underline{ }maria@ukr.net

\medskip
\textbf{Igor I.~Skrypnik}\\Institute of Applied Mathematics and Mechanics,
National Academy of Sciences of Ukraine, \\ \indent Heneral Batiuk Str. 19, 84116 Sloviansk, Ukraine\\
Vasyl' Stus Donetsk National University,
\\ \indent 600-richcha Str. 21, 21021 Vinnytsia, Ukraine\\ihor.skrypnik@gmail.com

\medskip
\textbf{Yevgeniia A. Yevgenieva}
\\ Max Planck Institute for Dynamics of Complex Technical Systems, \\ \indent Sandtorstrasse 1, 39106 Magdeburg, Germany
\\Institute of Applied Mathematics and Mechanics,
National Academy of Sciences of Ukraine, \\ \indent Heneral Batiuk Str. 19, 84116 Sloviansk, Ukraine\\yevgeniia.yevgenieva@gmail.com

\end{document}